# A new framework for nonlinear Kalman filters

### v6

@@ 6 The detailed algorithm of various nonlinear Kalman filters used in the paper
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
@@
11:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
12:     $\boldsymbol{\mathcal{X}}_{0,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}$
13:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
14:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
15:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=0,1,\ldots,2n_{x}$
16:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{X}}_{i,k|k-1}$
17:     $\boldsymbol{P}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}$
18:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
@@
20:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
21:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
22:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
24:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
25:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
26:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
27:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
28:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
29:    Recalibrate:
30:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
31:     $\boldsymbol{\mathcal{X}}_{0,k|k}=\boldsymbol{\hat{x}}_{k|k}$
32:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
33:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
34:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
35:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k}$
36:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
37:     $\boldsymbol{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
38:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
39:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
40:    Back out:
41:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
42:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
43:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
44:     end if
45:end for
Algorithm 6 The cubature Kalman filter with the new framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
8:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
9:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
10:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=1,\ldots,2n_{x}$
11:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
12:     $\boldsymbol{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
13:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
14:    Update:
15:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
16:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
17:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
18:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
19:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
20:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
21:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
22:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
24:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
25:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
26:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
27:    Recalibrate:
28:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
29:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
30:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
31:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k}$
32:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
33:     $\boldsymbol{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
34:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
35:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
36:    Back out:
37:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
38:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
39:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
40:     end if
41:end for
Acknowledgment

### v10

@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
@@
11:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
12:     $[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
13:     $\bm{\mathcal{X}}_{0,k-1}=\bm{\hat{x}}_{k-1|k-1}$
14:     $\bm{\mathcal{X}}_{i,k-1}=\bm{\hat{x}}_{k-1|k-1}+\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
15:     $\bm{\mathcal{X}}_{i+n_{x},k-1}=\bm{\hat{x}}_{k-1|k-1}-\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
16:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{f}(\bm{\mathcal{X}}_{i,k-1},\bm{u}_{k-1}),i=0,1,\ldots,2n_{x}$
17:     $\bm{\hat{x}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{X}}_{i,k|k-1}$
18:     $\bm{P}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
@@
20:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k})$
21:     $\bm{\mathcal{X}}_{0,k|k-1}=\bm{\hat{x}}_{k|k-1}$
22:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{\hat{x}}_{k|k-1}+\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
23:     $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
24:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
25:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k-1}$
26:     $\bm{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
27:     $\bm{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
28:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
29:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
30:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
31:    Recalibrate:
32:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1}),i=0,\ldots,2n_{x}$
33:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
34:     $\bm{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k}$
35:     $\bm{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
36:     $\bm{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
37:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
38:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
39:    Back out:
40:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
41:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
42:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
43:     end if
44:end for